Suppose that the Market for Cigarette is facing the Demand function Q = 20 – 2P and Supply function Q = 

10.5 + 0.5P:

a) What is the effect on the Equilibrium Price and Quantity when Government imposes a 7% of tax as 

percent of equilibrium price on each unit of Cigarette produced? [5 marks]

b) What is the price elasticity of demand at equilibrium after tax and comment on the answer? 

 [5 marks]


1
Expert's answer
2021-04-13T07:15:38-0400

Qd = 20 – 2P, Qs = 10.5 + 0.5P.

a) In equilibrium Qd = Qs, so:

20 – 2P = 10.5 + 0.5P,

P = 3.8,

Q = 20 - 2×3.8 = 12.4 units.

If government imposes a 7% of tax as percent of equilibrium price on each unit of Cigarette produced, then the equilibrium price will increase and the equilibrium quantity will decrease.

b) The price elasticity of demand will be higher at equilibrium after tax.